# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ Appendix E Training and Augmentation time
GMM-GDA 774.47 101.56 438.39
Appendix F Datasets and Implementation Details
F.1 Graph Classification
Characteristics and information about the datasets utilized for the graph classification task are presented in Table 8. As outlined in the main paper, we conduct experiments on IMDB-BINARY, IMDB-MULTI, PROTEINS, MUTAG, and DD, all sourced from the TUD Benchmark Ivanov et al. (2019). These datasets consist of either molecular or social graphs.
@@ F.1 Graph Classification
Dataset #Graphs Avg. Nodes Avg. Edges #Classes
IMDB-BINARY 1,000 19.77 96.53 2
IMDB-MULTI 1,500 13.00 65.94 3
MUTAG 188 17.93 19.79 2
PROTEINS 1,113 39.06 72.82 2
DD 1,178 284.32 715.66 2
F.2 Implementation Details
For all the used models, the same number of layers, hyperparameters, and activation functions were used. The models were trained using the cross-entropy loss function with the Adam optimizer, the number of epochs and learning rate were kept similar for the different approaches across all experiments. In Table 9, we present the optimal number of Gaussian distributions in the GMM for each dataset and GNN backbone
@@ F.2 Implementation Details
Dataset IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
@@
Appendix G Graph Distance Metrics
Let us consider the graph space $(\mathbb{G},\lVert\cdot\rVert_{\mathbb{G}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{G}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. When considering only structural changes, with fixed node features, the distance between two graphs $\mathcal{G}^{\lambda},\mathcal{G}$ is defined as
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\lVert A-A^{\lambda}\rVert_{\mathbb{G}},$ (8)
where $A^{\lambda},A$ are respectively the adjacency matrix of $\mathcal{G}^{\lambda},\mathcal{G},$ and the norm $\lVert\cdot\rVert_{\mathbb{G}}$ can be for example the Frobenius or spectral norm. If both structural and feature changes are considered, the distance extends to:
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\alpha\lVert A-A^{\lambda}\rVert_{\mathbb{G}}+\beta\lVert X-X^{\lambda}\rVert_{\mathbb{X}},$ (9)
where $X^{\lambda},X$ are the node feature matrices of $\mathcal{G}^{\lambda},\mathcal{G}$ respectively, and $\alpha,\beta$ are hyperparameters controlling the contribution of structural and feature differences.
In most baselines graph augmentation techniques, such as for instance $\mathcal{G}$-Mixup, SubMix, and DropNode, the alignment between nodes in the original graph $\mathcal{G}$ and the augmented graph $\mathcal{G}^{\lambda}$ is known. However, in cases where the node alignment is unknown, we must take into account node permutations. The distance between the two graphs is then defined as
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\min_{P\in\Pi}\left(\alpha\lVert A-PA^{\lambda}P^{T}\rVert_{\mathbb{G}}+\beta\lVert X-PX^{\lambda}\rVert_{\mathbb{X}}\right),\qquad\qquad$ (10)

### v3

@@ Appendix G Training and Augmentation time
GRATIN 774.47 101.56 438.39
Table 8: Classification accuracy ($\pm$ std) on different benchmark datasets for the data augmentation baselines using the GCN backbone. Higher accuracy (in %) is better.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
DropEdge $71.40{\scriptstyle\pm 3.69}$ $48.47{\scriptstyle\pm 2.36}$ $73.88{\scriptstyle\pm 7.98}$ $67.56{\scriptstyle\pm 4.50}$ $66.04{\scriptstyle\pm 4.35}$
DropNode $71.80{\scriptstyle\pm 4.11}$ $48.60{\scriptstyle\pm 3.45}$ $72.78{\scriptstyle\pm 8.01}$ $67.83{\scriptstyle\pm 4.75}$ $66.55{\scriptstyle\pm 3.97}$
SubMix $72.50{\scriptstyle\pm 3.90}$ $47.93{\scriptstyle\pm 3.58}$ $71.72{\scriptstyle\pm 10.59}$ $59.74{\scriptstyle\pm 2.83}$ $62.25{\scriptstyle\pm 3.29}$
Table 9: Classification accuracy ($\pm$ std) on different benchmark datasets for the data augmentation baselines using the GIN backbone. Higher accuracy (in %) is better.
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
DropEdge $71.70{\scriptstyle\pm 3.25}$ $43.67{\scriptstyle\pm 6.58}$ $72.36{\scriptstyle\pm 10.40}$ $66.31{\scriptstyle\pm 5.49}$ $70.20{\scriptstyle\pm 2.85}$
DropNode $71.00{\scriptstyle\pm 4.49}$ $42.67{\scriptstyle\pm 5.54}$ $75.00{\scriptstyle\pm 6.67}$ $64.60{\scriptstyle\pm 4.52}$ $70.80{\scriptstyle\pm 3.31}$
SubMix $71.10{\scriptstyle\pm 4.15}$ $42.80{\scriptstyle\pm 3.93}$ $84.53{\scriptstyle\pm 6.66}$ $60.10{\scriptstyle\pm 3.78}$ $70.03{\scriptstyle\pm 2.61}$
Furthermore, the performance of simple augmentation baselines such as DropEdge, DropNode, and SubMix significantly drops when using only one augmentation per training graph for all the epochs, which is the framework adopted in GRATIN, as shown in Tables 8 and 9. Tables 8 and 9, these baselines experience a significant drop in accuracy under this constraint, emphasizing their reliance on generating diverse augmentations at each epoch to maintain strong performance, c.f. the results in Tables 1 and 2.
The only graph augmentation baseline with comparable or better time complexity than GRATIN  is $\mathcal{G}$-Mixup. However, GRATIN  consistently outperforms $\mathcal{G}$-Mixup in most cases, as shown by the results in Tables 1 and 2.
Appendix H Augmentation Strategies
We present several graph data augmentation strategies, each defined by the augmentation function $A_{\lambda}$. For simplicity, we denote it $A_{\lambda}(\mathcal{G}_{n})$ instead of the more explicit $A_{\lambda}(\mathcal{G}_{n},y_{n})$.
Edge Perturbation. Randomly adding or removing edges. Given a graph $\mathcal{G}_{n}=(\mathcal{V}_{n},\mathcal{E}_{n},\mathbf{X}_{n})$, edge perturbation is defined as $A_{\lambda}(\mathcal{G}_{n})=(\mathcal{V}_{n},\mathcal{E}_{n}\cup\mathcal{E}_{\text{add}}\setminus\mathcal{E}_{\text{remove}},\mathbf{X}_{n}),$ where edges in $\mathcal{E}_{\text{add}}$ and $\mathcal{E}_{\text{remove}}$ are sampled from a Bernoulli distribution $\mathcal{P}(\lambda)=\mathcal{B}(\lambda)$ with probability $\lambda$.
Node Feature Perturbation. Augmenting node features by introducing noise or masking some features. This is given by $A_{\lambda}(\mathcal{G}_{n})=(\mathcal{V}_{n},\mathcal{E}_{n},\mathbf{X}_{n}+\lambda\mathbf{Z}),$ where $\mathbf{Z}\sim\mathcal{N}(0,I)$ for Gaussian noise addition.
Subgraph Sampling. Extracting subgraphs. A common approach is $k$-hop neighborhood sampling, where $A_{\lambda}(\mathcal{G}_{n})=(\mathcal{V}_{n}^{\prime},\mathcal{E}_{n}^{\prime},\mathbf{X}_{n}^{\prime}),$ with $\mathcal{V}_{n}^{\prime}\subseteq\mathcal{V}_{n}$ and $\mathcal{E}_{n}^{\prime}$ being edges induced by the $k$-hop neighbors of randomly selected nodes selected based on a prior distribution $\mathcal{P}(\lambda)$.
GRATIN. In our approach, the augmented hidden representations $\mathbf{h}_{\widetilde{\mathcal{G}}}=A_{\lambda}(\mathbf{h}_{\mathcal{G}})$ corresponds to a sampled vector from the GMM distribution $\mathcal{P}_{c}$ that was previously fit on the hidden representations $\mathcal{H}_{c}=\{\mathbf{h}_{\mathcal{G}}~{}~{}\mid y_{n}=c\}$ of the graphs in the training set with the same class $c$. Formally, $\mathbf{h}_{\widetilde{\mathcal{G}}}=A_{\lambda_{c}}(\{\mathcal{H}_{c}\mid\mathcal{G}\in c\})$, where $\lambda_{c}$ are the parameter of the GMM distribution $\mathcal{P}_{c}$.
It is important to note that in some augmentation strategies, the augmented graphs $\widetilde{\mathcal{G}}_{n}^{m}$ may not explicitly depend on the specific training graph $\mathcal{G}_{n}$. Instead, they may be sampled or generated based on other factors, such as a general graph distribution or global augmentation rules. This flexibility allows the augmentation framework to capture a broader range of variations while maintaining consistency with the original training data.
Appendix I Graph Distance Metrics
Let us consider the graph structure space $(\mathbb{A},\lVert\cdot\rVert_{\mathbb{A}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{G}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. When considering only structural changes with fixed node features, the distance between two graphs $\widetilde{\mathcal{G}},\mathcal{G}$ is defined as
$\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|=\lVert\mathbf{A}-\mathbf{\widetilde{A}}\rVert_{\mathbb{G}},$ (4)
where $\mathbf{\widetilde{A}},\mathbf{A}$ are respectively the adjacency matrix of $\widetilde{\mathcal{G}},\mathcal{G},$ and the norm $\lVert\cdot\rVert_{\mathbb{G}}$ can be for example the Frobenius or spectral norm. If both structural and feature changes are considered, the distance extends to:
$\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|=\alpha\lVert\mathbf{A}-\mathbf{\widetilde{A}}\rVert_{\mathbb{A}}+\beta\lVert\mathbf{X}-\widetilde{\mathbf{X}}\rVert_{\mathbb{X}},$ (5)
where $\widetilde{\mathbf{X}},\mathbf{X}$ are the node feature matrices of $\widetilde{\mathcal{G}},\mathcal{G}$ respectively, and $\alpha,\beta$ are hyperparameters controlling the contribution of structural and feature differences.
In most baseline graph augmentation techniques, such as $\mathcal{G}$-Mixup, SubMix, and DropNode, the alignment between nodes in the original graph $\mathcal{G}$ and the augmented graph $\widetilde{\mathcal{G}}$ is known. However, in cases where the node alignment is unknown, we must take into account node permutations. The distance between the two graphs is then defined as
$\left\|\widetilde{\mathcal{G}}-\mathcal{G}\right\|=\min_{P\in\Pi}\left(\alpha\lVert\mathbf{A}-P\mathbf{\widetilde{A}}P^{T}\rVert_{\mathbb{A}}+\beta\lVert\mathbf{X}-P\widetilde{\mathbf{X}}\rVert_{\mathbb{X}}\right),\qquad\qquad$ (6)
A common choice for measuring the distance between two graphs is the norm applied to their adjacency matrices. One widely used norm is the Frobenius norm, defined as $\lVert\mathbf{A}-\mathbf{\widetilde{A}}\rVert_{F}=\sqrt{\sum_{i,j}(A_{ij}-\widetilde{A}_{ij})^{2}}$, which captures element-wise differences between the adjacency matrices. Another commonly used norm is the spectral norm, defined as $\lVert\mathbf{A}-\mathbf{\widetilde{A}}\rVert_{2}=\sigma_{\max}(\mathbf{A}-\mathbf{\widetilde{A}})$, where $\sigma_{\max}$ denotes the largest singular value of the difference matrix.
Appendix J Gaussian Mixture Models
GMMs are probabilistic models used for modeling complex data by representing them as a mixture of multiple Gaussian distributions. The probability density function $p(\mathbf{x})$ of a data point $\mathbf{x}$ in a GMM with $K$ Gaussian components is given by:
$p(\mathbf{x}):=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}),$ (7)
where $\pi_{k}$ is the weight of the $k$-th Gaussian component, with $\pi_{k}\geq 0$ and $\sum_{k=1}^{K}\pi_{k}=1$, and $\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ is the Gaussian probability density function for the $k$-th component, defined as,
$\mathcal{N}(\mathbf{x}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k}):=\frac{1}{(2\pi)^{d/2}\det(\boldsymbol{\Sigma}_{k})^{1/2}}\times\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{k})^{\top}\boldsymbol{\Sigma}_{k}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{k})\right),$
where $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ are respectively the mean vectors and the covariance vectors of the $k$-th Gaussian component, and $d$ the dimensionality of $\mathbf{x}$. The parameters of a GMM are typically estimated using the EM algorithm (Dempster et al., 1977), which alternates between estimating the membership probabilities of data points for each Gaussian component (Expectation step) and updating the parameters of the Gaussian distributions (Maximization step). GMMs are a powerful tool in statistics and machine learning and are used for various purposes, including clustering and density estimation (Ozertem & Erdogmus, 2011; Naim & Gildea, 2012; Zhang et al., 2021).
Appendix K Experimental Setup
In this section, we detail the experimental setup for the conducted experiments. The necessary code to reproduce all our experiments is available on github at: https://github.com/abbahaddou/GRATIN
Datasets. We evaluate our model on five widely used datasets from the GNN literature, specifically IMDB-BIN, IMDB-MUL, PROTEINS, MUTAG, and DD, all sourced from the TUD Benchmark (Morris et al., 2020). These datasets consist of either molecular or social graphs. Detailed statistics for each dataset are provided in Table 10. We split the dataset into train/test/validation set by 80%/10%/10% and use 10-fold cross-validation for evaluation following the recent work of Zeng et al. (2024). If a dataset does not contain node features, we follow the standard practice in GNN literature by using one-hot encoding of node degrees as input features.
@@ Appendix I Graph Distance Metrics
Dataset #Graphs #Features Avg. Nodes Avg. Edges #Classes
IMDB-BIN 1,000 - 19.77 96.53 2
IMDB-MUL 1,500 - 13.00 65.94 3
MUTAG 188 7 17.93 19.79 2
PROTEINS 1,113 3 39.06 72.82 2
DD 1,178 82 284.32 715.66 2
COLLAB 5,000 - 74.5 4914.4 3
REDDIT-M5K 4,999 - 508.5 1189.7 5
Baselines. We benchmark the performance of our approach against the state-of-the-art graph data augmentation strategies. In particular, we consider the DropNode (You et al., 2020), DropEdge (Rong et al., 2020), SubMix (Yoo et al., 2022), $\mathcal{G}$-Mixup (Han et al., 2022) and GeoMix (Zeng et al., 2024). For the robustness to structure corruption experiment, c.f. Table 3, we also included NoisyGNN (Ennadir et al., 2024), a recent method explicitly designed to enhance GNN robustness, as an additional baseline.
Implementation Details. We used the PyTorch Geometric (PyG) open-source library, licensed under MIT (Fey & Lenssen, 2019). The experiments were conducted on an RTX A6000 GPU. For the datasets from the TUD Benchmark, we used a size base split. We utilized two GNN architectures, GIN and GCN, both consisting of two layers with a hidden dimension of 32. The GNN was trained on graph classification tasks for 300 epochs with a learning rate of $10^{-2}$ using the Adam optimizer (Kingma & Ba, 2014). To model the graph representations of each class, we fit a GMM using the EM algorithm, running for 100 iterations or until the average lower bound gain dropped below $10^{-3}$. The number of Gaussians used in the GMM is provided in Table 11. After generating new graph representations from each GMM, we fine-tuned the post-readout function for 100 epochs, maintaining the same learning rate of $10^{-2}$.
@@
Model IMDB-BIN IMDB-MUL MUTAG PROTEINS DD
@@
Appendix L Experiments on Large Datasets
We conducted additional experiments on larger-scale datasets, including COLLAB and REDDIT-MULTI-5K (Morris et al., 2020). The resufffor and GIN backbones are included in Table 12 and 13, which further confirm the effectiveness of our approach on graphs with a larger number of nodes and edges.
Table 12: Classification accuracy ($\pm$ std) on large datasets for the data augmentation baselines using the GCN backbone. Higher accuracy (in %) is better. The symbol ‘–’ denotes instances where the method has an excessive augmentation time, i.e., exceeding 2 hours
Vanilla DropEdge DropNode SubMix $\mathcal{G}-$Mixup GeoMix GRATIN
COLLAB $79.94\pm 1.61$ $79.70\pm 1.10$ $79.62\pm 1.84$ $81.86\pm 1.62$ $81.76\pm 1.58$ $80.74\pm 1.89$ $\mathbf{82.28\pm 1.82}$
REDDIT-MULTI-5K $48.88\pm 2.31$ $48.87\pm 1.99$ $48.73\pm 2.39$ $48.77\pm 2.01$ $46.23\pm 2.74$ $\mathbf{49.31\pm 1.56}$
Table 13: Classification accuracy ($\pm$ std) on large datasets for the data augmentation baselines using the GIN backbone. Higher accuracy (in %) is better. The symbol ‘–’ denotes instances where the method has an excessive augmentation time, i.e., exceeding 2 hours
Vanilla DropEdge DropNode SubMix $\mathcal{G}-$Mixup GeoMix GRATIN
COLLAB $77.80\pm 1.53$ $78.26\pm 1.46$ $78.86\pm 2.09$ $80.98\pm 1.24$ $78.89\pm 2.33$ $78.20\pm 1.31$ $\mathbf{79.08\pm 1.13}$
REDDIT-MULTI-5K $51.85\pm 4.29$ $44.52\pm 9.58$ $50.87\pm 3.36$ $49.93\pm 3.63$ $50.63\pm 4.04$ $\mathbf{51.53\pm 3.54}$
Appendix M Softmax Confidence and Entropy Distributions
One of the critical challenges in training GNN is Softmax saturation, where the model produces confident predictions. This high confidence leads to vanishing gradients, making the influence score of the augmented graphs converge to zero. To analyze this phenomenon, we examine the Softmax confidence and entropy distributions for GCN and GIN across the different graph classification datasets. In Figure 4, 5, 6, 7 and 8, we present histograms illustrating the distribution of maximum Softmax confidence and entropy for models trained on the original DD, IMDB-BIN, IMDB-MUL, MUTAG, and PROTEINS datasets. Each dataset panel contains two histograms, the Confidence Histogram and the Entropy Distribution. The Confidence Histogram illustrates the distribution of the maximum Softmax confidence scores assigned by the model to the predicted class. A higher confidence value indicates that the model is more certain about its classification decision. The Entropy Distribution provides a measure of uncertainty in the model’s predictions, computed as: $H(y)=-\sum_{c}y_{c}~{}\log(y_{c})$, where, for each class $c$, $y_{c}$ is the predicted probability corresponding to this class. Lower entropy values reflect high-confidence predictions and higher entropy values indicate more significant uncertainty.
Specifically, in the DD dataset, we observe an extreme case of Softmax saturation in GIN, where almost all predictions collapse to near-maximum confidence. The entropy histogram further reinforces the Softmax saturation in GIN on DD, where entropy values are heavily skewed towards zero, meaning the model rarely assigns significant probability mass outside the predicted class. Compared to other dataset-model settings, this effect is particularly pronounced in GIN trained on DD. In contrast, GCN exhibits a wider confidence distribution, maintaining a more balanced uncertainty.
Figure 4: Softmax confidence and entropy distributions for the IMDB-BIN dataset.
Figure 5: Softmax confidence and entropy distributions for the IMDB-MUL dataset.
Figure 6: Softmax confidence and entropy distributions for the MUTAG dataset.
Figure 7: Softmax confidence and entropy distributions for the PROTEINS dataset.
Figure 8: Softmax confidence and entropy distributions for the DD dataset.